OpenStudy (anonymous):

Medal PLEASE Help Which is an example of the commutative property? A. 21 + (–21) = 0 B. 13 + (–19) = –19 + 13 C. –(12 + 7) = –12 + (–7) D. 11 + (–3 + 8) = (11 + –3) + 8

OpenStudy (anonymous):

The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2. So the anser would be....

OpenStudy (anonymous):

A+B = B+A

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Or B?

OpenStudy (anonymous):

no the answer is D because your have to flip the equation.

OpenStudy (anonymous):

Do you get it now? 32+12=12+32